# Fast Tensor Completion via Approximate Richardson Iteration

### v1

@@ 3 Approximate Richardson Iteration
We now present our main techniques for reducing tensor completion to tensor decomposition. When we use ALS to solve a TC problem, we must efficiently solve least-squares problems $\min_{\mathbf{x}}\,\lVert\mathbf{Px}-\mathbf{q}\rVert_{2}$. The rows of the design matrix $\mathbf{P}$ correspond to the subset of observations in the TC problem, so $\mathbf{P}$ does not necessarily have the structure of the design matrix in the full TD problem.
@@
We show that our novel approximate-mini-ALS method is a principled approach for tensor completion. In Section 3.1, we show how lifting restores the structure of the full TD ALS update step, enabling fast least-squares methods for a larger (but equivalent) problem. In Section 3.2, we show that iteratively solving this lifted problem (i.e., mini-ALS) is connected to an iterative method called the Richardson iteration (Richardson, 1911), which we can also view as a matrix-splitting method. In other words, mini-ALS and the Richardson iteration with a certain preconditioner give the same sequence of iterates $\{\mathbf{x}^{(k)}\}_{k\geq 0}$. Lastly in Section 3.3, we prove novel convergence guarantees for approximately solving the lifted problem (i.e., for approximate-mini-ALS). This allows us to directly use fast leverage-score sampling algorithms for CP decomposition (Cheng et al., 2016; Larsen and Kolda, 2022; Bharadwaj et al., 2023), Tucker decomposition (Diao et al., 2019; Fahrbach et al., 2022), and TT decomposition (Bharadwaj et al., 2024) as blackbox subroutines.
@@ 3.1 Lifting to a Structured Problem
$\mathbf{x}^{*}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{P}\mathbf{x}-\mathbf{q}\rVert_{2}^{2}\,.$ (7)
@@
then $\mathbf{x}^{*}$ also minimizes (7), i.e., the original linear regression problem $\min_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{P}\mathbf{x}-\mathbf{q}\rVert_{2}^{2}$.
@@
Lemma 3.2.
Problem 8 is a convex quadratic program.
@@
Remark 3.3.
Problem 8 is not a linear regression problem with (structured) design matrix $\mathbf{A}$ since there are $\mathbf{b}_{\overline{\Omega}}$ variables in the response. There is, however, enough structure to employ block minimization to alternate between minimizing $\mathbf{x}$ and $\mathbf{b}_{\overline{\Omega}}$.
3.2 Iterative Methods for the Lifted Problem
Iterative methods for solving linear systems and regression problems have a long history and have been used to expedite several algorithms in theory and practice. The algorithms we consider use the exact arithmetic model, but all of these methods can be carried out with numbers with $\log\nicefrac{{\kappa}}{{\varepsilon}}$ bits, where $\kappa$ is the condition number of the matrix (see, e.g., Ghadiri et al. (2023b, 2024)). There is a literature on inexact Richardson iteration for solving linear systems, but they require the error $\widehat{\varepsilon}$ to be smaller than than $1/\kappa$, which is not achievable with leverage-score sampling Golub and Overton (1988); Golub and van der Vorst (1997).
Lemma 3.4 (Preconditioned Richardson iteration, (Lee and Vempala, 2024, Lemma 6.1)).
Then, we have that
$\lVert\mathbf{x}^{(k+1)}-\mathbf{x}^{*}\rVert_{\mathbf{M}}\leq\Bigl{(}1-\frac{1}{\beta}\Bigr{)}\lVert\mathbf{x}^{(k)}-\mathbf{x}^{*}\rVert_{\mathbf{M}}\,.$
Lemma 3.5.
$\mathbf{x}^{(k+1)}=\mathbf{x}^{(k)}-(\mathbf{A}^{\top}\mathbf{A})^{-1}(\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{q}})\,.$ (9)
@@ 3.2 Iterative Methods for the Lifted Problem
$=\mathbf{x}^{(k)}-(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}(\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{q}})\,.$
@@
$\mathbf{A}^{\top}(\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{q}})$ $=\bigl{(}\widetilde{\mathbf{P}}-(\widetilde{\mathbf{P}}-\mathbf{A})\bigr{)}^{\top}\bigl{[}\begin{matrix}\widetilde{\mathbf{P}}&\widetilde{\mathbf{q}}\end{matrix}\bigr{]}\begin{bmatrix}\mathbf{x}^{(k)}\\ -1\end{bmatrix}$
$=\widetilde{\mathbf{P}}^{\top}(\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{q}})\,.$
@@
$\mathbf{x}^{(k+1)}=\mathbf{x}^{(k)}-(\mathbf{A}^{\top}\mathbf{A})^{-1}(\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{q}})\,,$
@@
Remark 3.6.
$\mathbf{x}^{(k+1)}$ $=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{Ax}-\widetilde{\mathbf{q}}^{(k)}\rVert^{2}_{2}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\bigl{\{}\lVert\mathbf{Px}-\mathbf{q}\rVert^{2}_{2}+\lVert\mathbf{A}_{\overline{\Omega}}\,(\mathbf{x}-\mathbf{x}^{(k)})\rVert_{2}^{2}\bigr{\}}\,.$
3.3 Approximately Solving the Lifted Problem
We have shown that alternating minimization for the lifted problem (8) has strong connections to preconditioned Richardson iteration and inherits its convergence guarantees. However, for this observation to be useful, we need to use fast regression algorithms for the $\mathbf{x}^{(k+1)}$ updates that exploit the structure of $\mathbf{A}$, i.e., when solving $\min_{\mathbf{x}}\,\lVert\mathbf{Ax}-\widetilde{\mathbf{q}}^{(k)}\rVert_{2}$.
This is where leverage score sampling comes in to play. We exploit the structure of $\mathbf{A}$ to efficiently compute its leverage scores, and then we solve the regression problem efficiently but approximately. By using a sketching method, our work deviates from the standard (exact) Richardson iteration.
Result: $\widetilde{\mathbf{x}}\in\mathbb{R}^{R}$
1 Initialize $\mathbf{x}^{(0)}=\bm{0}$
2
return $\mathbf{x}^{(k)}$
Algorithm 1 approx-mini-als
Theorem 3.7.
$\lVert\widetilde{\mathbf{P}}\widetilde{\mathbf{x}}-\widetilde{\mathbf{q}}\rVert_{2}^{2}$ $\leq\left(1+\frac{2\widehat{\varepsilon}}{(\nicefrac{{1}}{{\beta}}-\sqrt{\widehat{\varepsilon}})^{2}}\right)\,\min_{\mathbf{x}}\,\lVert\widetilde{\mathbf{P}}\mathbf{x}-\widetilde{\mathbf{q}}\rVert_{2}^{2}+\varepsilon\,\lVert\widetilde{\mathbf{P}}\,(\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{P}})^{-1}\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{q}}\rVert_{2}^{2}\,,$
@@ 1 Initialize $\mathbf{x}^{(0)}=\bm{0}$
$=\mathbf{x}^{(k)}-(\mathbf{A}^{\top}\mathbf{A})^{-1}(\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{P}}\mathbf{x}^{(k)}-\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{q}})\,,$

### v2

@@ 3 Approximate Richardson Iteration
We now present our main techniques for reducing tensor completion to tensor decomposition. When using ALS to solve a TC problem, we must efficiently solve least-squares problems
$\min_{\mathbf{x}}\,\lVert\mathbf{Px}-\mathbf{q}\rVert_{2}.$
The rows of the design matrix $\mathbf{P}$ correspond to the subset of observations in the TC problem, which means $\mathbf{P}$ does not necessarily have the structure of the design matrix in the full TD problem.
@@
We show that our approximate-mini-ALS method is a principled approach for tensor completion. In Section 3.1, we prove that lifting restores the structure of the full TD ALS update step, enabling fast least-squares methods for a larger but equivalent problem. In Section 3.2, we show that iteratively solving the lifted problem (i.e., mini-ALS) is connected to an iterative method called the Richardson iteration (Richardson, 1911), which we can also view as a matrix-splitting method. In other words, mini-ALS and the Richardson iteration with a certain preconditioner give the same sequence of iterates $\{\mathbf{x}^{(k)}\}_{k\geq 0}$. Lastly in Section 3.3, we prove novel convergence guarantees for approximately solving the lifted problem (i.e., for approximate-mini-ALS). This allows us to directly use fast leverage-score sampling algorithms for CP decomposition (Cheng et al., 2016; Larsen & Kolda, 2022; Bharadwaj et al., 2023), Tucker decomposition (Diao et al., 2019; Fahrbach et al., 2022), and TT decomposition (Bharadwaj et al., 2024) as blackbox subroutines. All missing proofs are deferred to Appendix B.
@@ 3.1 Lifting to a Structured Problem
$\mathbf{x}^{*}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{P}\mathbf{x}-\mathbf{q}\rVert_{2}\,.$ (7)
@@
then $\mathbf{x}^{*}$ also minimizes (7), i.e., the original linear regression problem $\min_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{P}\mathbf{x}-\mathbf{q}\rVert_{2}$.
Lemma 3.2.
Problem 8 is a convex quadratic program.
Remark 3.3.
Problem 8 is not a linear regression problem with (structured) design matrix $\mathbf{A}$ because there are $\mathbf{b}_{\overline{\Omega}}$ variables in the response. However, there is enough structure to employ block minimization to alternate between minimizing $\mathbf{x}$ and $\mathbf{b}_{\overline{\Omega}}$.
3.2 Iterative Methods for the Lifted Problem
Iterative methods for solving linear systems and regression problems have a long history and have been used to speed up several algorithms in theory and practice. The algorithms we consider use the exact arithmetic model, but all of these methods can be carried out with numbers with $\log\nicefrac{{\kappa}}{{\varepsilon}}$ bits, where $\kappa$ is the condition number of the matrix (see, e.g., Ghadiri et al. (2023b, 2024)). There is a literature on inexact Richardson iteration for solving linear systems, but they require the error $\widehat{\varepsilon}$ to be smaller than than $1/\kappa$, which is not achievable with leverage-score sampling (Golub & Overton, 1988; Golub & van der Vorst, 1997).
Lemma 3.4 (Preconditioned Richardson iteration, (Lee & Vempala, 2024, Lemma 6.1)).
Then, we have that
$\lVert\mathbf{x}^{(k+1)}-\mathbf{x}^{*}\rVert_{\mathbf{M}}\leq\left(1-\frac{1}{\beta}\right)\lVert\mathbf{x}^{(k)}-\mathbf{x}^{*}\rVert_{\mathbf{M}}\,.$
Lemma 3.5.
[…]
Remark 3.6.
$\mathbf{x}^{(k+1)}$ $=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\lVert\mathbf{Ax}-\widetilde{\mathbf{q}}^{(k)}\rVert^{2}_{2}$
$=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{R}}\,\bigl{\{}\lVert\mathbf{Px}-\mathbf{q}\rVert^{2}_{2}+\lVert\mathbf{A}_{\overline{\Omega}}\,(\mathbf{x}-\mathbf{x}^{(k)})\rVert_{2}^{2}\bigr{\}}\,.$
3.3 Approximately Solving the Lifted Problem
We have shown that alternating minimization for the lifted problem (8) is closely connected to preconditioned Richardson iteration and inherits its convergence guarantees. For this to be useful, we need to use fast regression algorithms for the $\mathbf{x}^{(k+1)}$ updates that exploit the structure of $\mathbf{A}$, i.e., when solving $\min_{\mathbf{x}}\,\lVert\mathbf{Ax}-\widetilde{\mathbf{q}}^{(k)}\rVert_{2}$, where $\mathbf{x}^{(k+1)}$ is the vector produced in iteration $k$ of Algorithm 1.
This is where leverage score sampling comes in to play. We exploit the structure of $\mathbf{A}$ to efficiently compute its leverage scores, and then we solve the regression problem efficiently but approximately.
Theorem 3.7.
$\lVert\widetilde{\mathbf{P}}\widetilde{\mathbf{x}}-\widetilde{\mathbf{q}}\rVert_{2}^{2}$ $\leq\left(1+\frac{2\widehat{\varepsilon}}{(\nicefrac{{1}}{{\beta}}-\sqrt{\widehat{\varepsilon}})^{2}}\right)\,\min_{\mathbf{x}}\,\lVert\widetilde{\mathbf{P}}\mathbf{x}-\widetilde{\mathbf{q}}\rVert_{2}^{2}$
$\quad+\varepsilon\,\lVert\widetilde{\mathbf{P}}\,(\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{P}})^{-1}\widetilde{\mathbf{P}}^{\top}\widetilde{\mathbf{q}}\rVert_{2}^{2}\,$,
Remark 3.8.
[…]
Result: $\widetilde{\mathbf{x}}\in\mathbb{R}^{R}$
1 Initialize $\mathbf{x}^{(0)}=\bm{0}$
2
return $\mathbf{x}^{(k)}$
Algorithm 1 approx-mini-als
Bounding $\beta$.
4 Sampling Methods for Tensor Completion
[…]
where $\mathbf{\pi}_{\mathbf{P}}$ and $\mathbf{\pi}_{\mathbf{P}^{\perp}}$ are the orthogonal projection matrices into the column space and null space of $\mathbf{P}$, respectively (see Section A.1). With this in hand, we are ready to present the running times of our lifted iterative method for TC problems by combining Theorem 3.7 with state-of-the-art tensor decomposition results based on leverage score sampling.
4.1 CP Completion
The design matrix $(\mathbf{A}^{\neq k})_{\Omega}$ does not necessarily have any structure, so a direct method relies on solving the normal equation, which takes $O(R^{\omega}+R|\Omega|(R+I_{k}))$ time. Thus, the running time of one round of ALS, i.e., updating all $N$ factors, is $O(N(R^{\omega}+R^{2}|\Omega|)+R|\Omega|\sum_{n=1}^{N}I_{n})$.
[…]
Corollary 4.1.
Note that there is no dependence on $|\Omega|$ in the running time due to leverage score sampling, i.e., it runs in sublinear time.
4.2 Tucker Completion
4.2.1 Core tensor update
Recall that for a Tucker decomposition we use the notation $I:=\prod_{n\in[N]}I_{n}$ and $R:=\prod_{n\in[N]}R_{n}$. The ALS core tensor update in the Tucker completion problem is
$\bm{\mathscr{G}}\leftarrow\operatorname*{arg\,min}_{\bm{\mathscr{G}}^{\prime}\in\mathbb{R}^{R_{1}\times\cdots\times R_{N}}}\left\lVert\left(\bigotimes_{n=1}^{N}\mathbf{A}^{(n)}\right)_{\Omega}\hskip-5.69046pt\operatorname*{vec}(\bm{\mathscr{G}}^{\prime})-\operatorname*{vec}(\bm{\mathscr{X}})_{\Omega}\right\rVert_{2}.$
Corollary 4.2.
$\tilde{O}\left(\left(\mathsf{C}+\frac{\beta^{2}R^{2-\theta^{*}}}{\varepsilon_{1}}\right)\beta\log\frac{1}{\varepsilon_{2}}\right)\,$, (11)
where $\mathsf{C}:=\sum_{n=1}^{N}(I_{n}R_{n}+\beta^{4}R_{n}^{\omega}N^{2}\varepsilon_{1}^{-2})$.
4.2.2 Factor matrix update
$\mathbf{A}^{(k)}\leftarrow\hskip-2.84544pt\operatorname*{arg\,min}_{\mathbf{A}\in\mathbb{R}^{I_{k}\times R_{k}}}\left\lVert\left(\left(\bigotimes_{n=1,n\neq k}^{N}\hskip-2.84544pt\mathbf{A}^{(n)}\right)\mathbf{G}_{(k)}^{\top}\right)_{\Omega}\hskip-2.84544pt\mathbf{A}^{\top}\hskip-2.84544pt-\mathbf{Q}\right\rVert_{\textnormal{F}},$
Corollary 4.3.
$\tilde{O}\left(\left(\mathsf{C}+\frac{\beta^{2}I_{k}R_{\neq k}^{2-\theta^{*}}}{\varepsilon_{1}}+I_{k}R\sum_{n=1}^{N}R_{n}\right)\beta\log\frac{1}{\varepsilon_{2}}\right)\,$,
where $\mathsf{C}=\sum_{n=1}^{N}(I_{n}R_{n}+\beta^{4}R_{n}^{\omega}N^{2}\varepsilon_{1}^{-2})$.
4.3 TT Completion
[…]
Solving this directly with the normal equation takes $O(\bar{R}_{k}^{\omega}+\bar{R}_{k}|\Omega|(\bar{R}_{k}+I_{k}))$ time for $\bar{R}_{k}:=R_{k-1}R_{k}$. Thus, the time for one round of ALS is $O(\sum_{n=1}^{N}(\bar{R}_{n}^{\omega}+\bar{R}_{n}|\Omega|(\bar{R}_{n}+I_{n})))$.
In contrast, Bharadwaj et al. (2024, Corollary 4.4) show that one round of approximate TT-core updates, if $R_{n}=R$ for all $n\in[N-1]$, can run in time $\tilde{O}(R^{4}\varepsilon^{-1}\sum_{n=1}^{N}\left(N+I_{n}\right))$. See Appendix C for technical details.
Corollary 4.4.
5 Experiments
In this section, we study the empirical performance of our algorithm and compare it to direct and expectation maximization (EM) methods for a coupled matrix problem and CP completion tasks. [^1] The code is available at https://github.com/fahrbach/fast-tensor-completion.
5.1 Warm-Up: Coupled Matrix Problem
First we consider the coupled matrix problem
where $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}\in\mathbb{R}^{n\times d}$ are given, $\mathbf{X},\mathbf{Y}\in\mathbb{R}^{d\times d}$ are unknown, and $\mathbf{E}\in\mathbb{R}^{n\times n}$ is a matrix with half of its entries randomly revealed (Baksalary & Kala, 1980). For fixed $\mathbf{Y}$, we compute $\mathbf{X}$ by solving the Kronecker regression problem
The matrix $\mathbf{Y}$ can be updated by solving a similar regression problem. Therefore, we can apply an alternating minimization algorithm to compute $\mathbf{X}$ and $\mathbf{Y}$. For these experiments, we initialize $\mathbf{X}=\mathbf{Y}=\mathbf{I}$. We present the results in Figure 1, which are averaged over five trials.
Data generation.
The entries of $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D},\mathbf{X},\mathbf{Y}$ are sampled independently from a uniform distribution on $[0,1)$, and then we set $\mathbf{E}\leftarrow\mathbf{A}\mathbf{X}\mathbf{B}^{\top}+\mathbf{C}\mathbf{Y}\mathbf{D}^{\top}$. We consider the setting where half of the entries of $\mathbf{E}$ are observed (chosen uniformly at random). We set $n=2000$ and $d=10$. Note that since we observe a subset of entries of $\mathbf{E}$, the Kronecker regression structure is lost.
Algorithms.
We compare the direct, mini-als, and approximate-mini-als methods. The latter two use adaptive step sizes based on the trajectory of the iterates $\mathbf{x}^{(k)}$ (see Section D.1.2 for details). direct solves the normal equation in each ALS step and runs in $O(|\Omega|R^{2}+R^{3})$ time since it computes $(\mathbf{P}^{\top}\mathbf{P})^{-1}$. mini-als is Algorithm 1 with $\widehat{\varepsilon}=0$ and $\varepsilon>0$. mini-als uses the Kronecker product properties $((\mathbf{B}\otimes\mathbf{A})^{\top}(\mathbf{B}\otimes\mathbf{A}))^{-1}=(\mathbf{B}^{\top}\mathbf{B})^{-1}\otimes(\mathbf{A}^{\top}\mathbf{A})^{-1}$ and $(\mathbf{B}\otimes\mathbf{A})\operatorname*{vec}(\mathbf{X})=\operatorname*{vec}(\mathbf{A}\mathbf{X}\mathbf{B}^{\top})$ for improved efficiency. approximate-mini-als uses leverage score sampling for Kronecker products similar to Fahrbach et al. (2022); Diao et al. (2019), which is a direct application of Corollary 4.2. For leverage score sampling, we sample $1\%$ of rows in each iteration of approximate-mini-als.
Figure 1: Coupled matrix results for $\mathbf{E}\in\mathbb{R}^{n\times n}$, $\mathbf{X},\mathbf{Y}\in\mathbb{R}^{d\times d}$ with $n=2000$ and $d=10$ that compare the direct method, mini-ALS, and approximate-mini-ALS via leverage score sampling.
Results.
The left plot in Figure 1 shows the total running time of these three algorithms across all ALS iterations. The right plots shows the mean squared error (MSE) at each step of ALS. Since approximate-mini-als is stochastic, it does not necessarily attain the minimal error in the matrix’s kernel space. i.e., the first term on the right‐hand side of (10) can exceed the minimum error by a factor of $(1+\varepsilon)$, which allows it to follow a different convergence path and achieve a lower MSE than the other methods around step 15 of ALS.
5.2 CP Completion
Now we compare methods for the CP completion task. For a given tensor $\bm{\mathscr{X}}$ and sample ratio $p\in[0,1]$, let $\bm{\mathscr{X}}_{\Omega}$ be a partially observed tensor with a random $p$ fraction of entries revealed. We fit $\bm{\mathscr{X}}_{\Omega}$ with a rank-$R$ CP decomposition by minimizing the training relative reconstruction error (RRE) $\lVert(\widehat{\bm{\mathscr{X}}}-\bm{\mathscr{X}})_{\Omega}\rVert_{\textnormal{F}}/\lVert\bm{\mathscr{X}}_{\Omega}\rVert_{\textnormal{F}}$ using different ALS algorithms.
Figure 2: Algorithm comparison for a low-rank CP completion task on the cardiac-mri and hyperspectral tensor datasets. The first column illustrates convergence rates and relative reconstruction error (RRE) of the ALS algorithm for various CP‐decomposition ranks. The second column plots RRE over ALS steps for direct, parafac, and our method (mini-als) under different $\varepsilon$ values. The third column shows total running time of these algorithms at varying sample ratios, i.e., number of observed entries. Finally, the fourth column displays total running times for direct, parafac, and accelerated-min-als (instead of mini-als) across different $\varepsilon$ settings.
Datasets.
We consider two real-world tensors. cardiac-mri is a $256\times 256\times 14\times 20$ tensor of MRI measurements indexed by $(x,y,z,t)$ where $(x,y,z)$ is a point in space and $t$ corresponds to time. hyperspectral is $1024\times 1344\times 33$ tensor of time-lapse hyperspectral radiance images (Nascimento et al., 2016). We also consider synthetic low-rank CP and Tucker tensors in Section D.1.1.
Algorithms.
We compare direct, parafac, mini-als, and accelerated-mini-als methods. direct solves the normal equation in each ALS step for the original problem of the form (4) and has running time $O(|\Omega|R^{2}+R^{3})$ since it computes $(\mathbf{A}_{\Omega}^{\top}\mathbf{A}_{\Omega})^{-1}$. mini-als is our lifting approach in Algorithm 1 with $\widehat{\varepsilon}=0$ and $\varepsilon>0$. parafac is the EM algorithm of Tomasi & Bro (2005), which is equivalent to running Algorithm 1 for exactly one iteration in each step of the ALS algorithm. accelerated-mini-als uses adaptive step sizes based on the trajectory of the iterates $\mathbf{x}^{(k)}$.
Results.
  • In the first column, we set $p=0.1$ and sweep over the rank $R$ using direct to demonstrate how the train RRE (solid line) and test RRE $\lVert\widehat{\bm{\mathscr{X}}}-\bm{\mathscr{X}}\rVert_{\textnormal{F}}/\lVert\bm{\mathscr{X}}\rVert_{\textnormal{F}}$ (dashed line) decrease as a function of $R$ and the ALS step count. Note that the test RRE is the loss on the entire (mostly unobserved) tensor, and is slightly above the train RRE curves at all times.
  • In the second column, we fix the parameters $(p,R)=(0.1,16)$ and study the solution quality of mini-als for different values of $\varepsilon$ compared to direct and parafac. As $\varepsilon$ decreases, we recover solutions in each step of ALS that are as good as direct, which agrees with the claim that lifting simulates the preconditioned Richardson iteration (Lemma 3.5).
  • In the fourth column, we compare the total running time of the accelerated-mini-als algorithm (dash-dot lines) to mini-als (third column). Our accelerated method extrapolates the trajectory of the iterates $\mathbf{x}^{(k)}$ during mini-ALS using a geometric series, which allows us to solve collinear iterates in one step. We illustrate this idea with Figure 4 in Section D.1.2. accelerated-mini-als achieves better solution quality than mini-als for a given $\varepsilon$ and always runs faster, especially for small values of $p$.
6 Conclusion
This work introduces a novel lifting approach for tensor completion. We build on fast sketching-based TD algorithms as ALS subroutines, extending their guarantees to the TC setting and establishing novel connections to iterative methods. We prove guarantees for the convergence rate of an approximate version of the Richardson iteration, and we study how these algorithms perform in practice on real-world tensors. One interesting future direction is to analyze the speedup that adaptive step sizes give in accelerated-mini-als.
Acknowledgements
We would like to thank Arkadi Nemirovski and Devavrat Shah for insightful discussions. MG and AJ acknowledge the support of ARO MURI W911-NF-19-1-0217 and ONR N00014-23-1-2299. YK is supported in part by NSF Award CCF-2106444, and did this work as a student researcher at Google Research.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential consequences of our work, but none that we feel must be highlighted here.
Appendix A Missing Details for Section 2
A.1 Least-Squares Linear Regression
A.2 Leverage Score Sampling for Tensor Decomposition
A.3 Tensor-Train Decomposition
[…]
A.4 Tensor Networks
A notable special case is a fully-connected tensor network (FCTN) (Zheng et al., 2021) decomposition. It consists of $N$ tensors of the same order, where in its tensor diagram, all pairs of nodes are connected as its name suggests.
ALS for TN decomposition.
[…]
Other tensor decompositions.
We also discuss other important special cases of TN decompositions including the tensor ring (TR) (Zhao et al., 2016) and tensor wheel (TW) (Wu et al., 2022) decompositions. Both decompositions can be succinctly described through tensor diagram notation.
A TR decomposition consists of $N$ tensors $\bm{\mathscr{A}}^{(1)},\dots,\bm{\mathscr{A}}^{(N)}$, where for each $k\in[N]$, the node of $\bm{\mathscr{A}}^{(k)}$ is connected to the two neighboring nodes of $\bm{\mathscr{A}}^{(k-1)}$ and $\bm{\mathscr{A}}^{(k+1)}$ (precisely, the superscripts here refer to modulo with respect to $N$) so that the resulting tensor diagram looks exactly like a ring. Since a fast TR decomposition through leverage-score sampling is facilitated by Malik & Becker (2021), our approach of lifting and completion can be straightforwardly applied to the TR decomposition.
A TW decomposition consists of one core tensor $\bm{\mathscr{C}}$ and $N$ factor tensors $\bm{\mathscr{A}}^{(1)},\dots,\bm{\mathscr{A}}^{(N)}$. As its name suggests, its tensor diagram is essentially that of the tensor ring (consisting of $N$ factor tensors) with the node representing $\bm{\mathscr{C}}$ connected to all factor tensors. Thanks to a recent work of Wang et al. (2024) using leverage-score sampling for the TW decomposition, our approach can be readily applied as well.
The t-SVD (Kilmer & Martin, 2011) and t-CUR (Tarzanagh & Michailidis, 2018) decompositions for third-order tensors that have been applied to spatiotemporal data. They factor a tensor via the t-product into three tensors; in the case of t-SVD, one of these factors is f-diagonal (see Kilmer & Martin (2011) for definitions). Tarzanagh & Michailidis (2018) propose fast, sampling-based algorithms for t-product–based tensor decompositions. These algorithms can be combined with our lifting method and thus extended to the tensor completion setting.
Appendix B Missing Analysis for Section 3
B.1 Proof of Lemma 3.1
See 3.1
@@ Appendix B Missing Analysis for Section 3
B.2 Proof of Lemma 3.2
See 3.2
@@
B.3 Proof of Lemma 3.5
See 3.5
@@
[…]
@@
[…]
@@
[…]
@@
B.4 Proof of Theorem 3.7
See 3.7
@@
[…]